OpenStudy (anonymous):

Rewrite in simplest rational exponent form √x • 4√x. Show each step of your process.

OpenStudy (jdoe0001):

\(\bf \sqrt{x}\cdot 4\sqrt{x}\implies 4\sqrt{x}\cdot \sqrt{x}\implies 4(\sqrt{x})^2\implies 4\sqrt{(x)^2}\implies ?\)

Nnesha (nnesha):

is it 4 or 4th root ?

Nnesha (nnesha):

oopz nvm .-.

OpenStudy (triciaal):